To save money for a new house, you want to begin contributing money to a savings account. Your plan is to make equal contributions to the account for eight years. Each contribution will be R18490,00. The contributions will come at the end of every three month period, starting three months from now. The savings account pays 9% interest per annum, compounded quarterly. How much money do you expect to have in the account at the end of the eight years?

1
Expert's answer
2022-02-27T11:46:21-0500

"FV=\frac{PV ((1+\frac{r}{m})^{nm}-1)}{(\frac{r}{m})}"


PV=R18490,00

r=9% or 0.09

m=4

n=8


"FV=\frac{R18490,00((1+\frac{0.09}{4})^{8×4}-1)}{(\frac{0.09}{4})}"


"FV=R853090"